# Wasserstein Convergence of Score-based Generative Models under Semiconvexity and Discontinuous Gradients

### v1

@@ Acknowledgments.
This work was supported by Innovate UK [grant number 10081810]. This work has received funding from the Ministry of Trade, Industry and Energy (MOTIE) and Korea Institute for Advancement of Technology (KIAT) through the International Cooperative R&D program (No.P0025828).
@@ Appendix A Regularity of the score function
Proposition A.1.
[15, Proposition 3.1] Let $\pi_{\mathsf{D}}$ be absolutely continuous with respect to the Lebesgue measure, and denote its density by $p_{0}$. The map defined in (34) is positive and solution of the following Fokker–Planck equation on $(0,T]\times\mathbb{R}^{d}$:
@@ Appendix B Further Details on Assumption 2 and Weak Convexity of the Data Distribution
Proof of Proposition 3.7.
We begin by considering that $\pi_{\mathsf{D}}$ satisfies Assumption 2. Recall that $f_{L}$ is defined as in (13). Note that $r\mapsto r^{-1}f_{L}(r)$ is non-increasing on $(0,\infty)$ and $f_{L}^{{}^{\prime}}(0)=L>r^{-1}f_{L}(r)$ for $r\in(0,R]$. We look for $L>0$ satisfying
@@
so as $L>K+\mu$. Note that $\tanh(x)\leq x$ for all $x\geq 0$ Therefore, if we choose $x=\frac{\sqrt{K+\mu}}{2}R$, then
@@
Using (37) and $\lim_{x\uparrow\infty}x\tanh(x)=\infty$, we deduce that there exists $x^{\star}>0$ such that
@@
with $x^{\star}>\frac{\sqrt{K+\mu}}{2}R$, since $x\mapsto x\tanh(x)$ is non-decreasing on $(0,\infty)$. By Assumption 2 and (35), we have
@@
where it is used that $r^{-1}f_{L}(r)>0$ for all $r>0$. This proves the first part of the statement in Proposition 3.7, i.e. the lower bound (16).
Conversely, assume that $U$ is weakly convex as in Definition 3.5 with lower bound (16) for some known constants $\mu$ and $L>0$. We look for $R$ such that
@@
where it is used that $r^{-1}f_{L}(r)$ is decreasing on $(0,\infty)$. Let $\widetilde{\mu}:=\mu-R^{-1}f_{L}(R)$, so (40) becomes $\kappa_{U}(r)\geq\widetilde{\mu}>0$, for all $r>R$. One notes that
@@
If $\mu>L$, (41) is satisfied for all $R>0$. If $\mu\leq L$, (41) holds for $R\geq R_{0}$, where $R_{0}$ is the unique solution to
@@
Since $\frac{\tanh(z)}{z}$ monotonically decreases from $1$ to 00 as $z$ increases, a unique $z_{0}>0$ solving (43) exists for $\mu<L$. Therefore, (40) is satisfied for $R\geq R_{0}=\frac{2z_{0}}{L^{1/2}}$. This proves that $U$ is $\widetilde{\mu}$-strongly convex at infinity, and therefore (17). Using the assumption that $U$ is weakly convex as in Definition 3.5, one obtains that
@@
We distinguish two cases for the lower bound in (44). If $\mu>L$, then $\kappa_{U}(r)\geq-K$ for $r\leq R$ for all $R>0$ and $K\geq 0$. If $\mu\leq L$, then, by setting $K=L-\mu$ in (43), we have $\kappa_{U}(r)\geq-K$ for $r\leq R$ for all $R>0$. This proves that $U$ is $K$-semiconvex, and therefore (18). This concludes the proof for the second part of the statement in Proposition 3.7.
Proof of Proposition 3.10.
@@
Note that (46) is satisfied for all $x>0$ when $K=0$. In addition, we have
@@
By (48), the function $g$ in (46) has a minimum at $\frac{K}{\mu^{2}}$ and
@@
for all $K,\mu>0$. By (47) and (48), there exists $x>\frac{K}{\mu^{2}}$ such that (46) is strictly positive. Therefore, there exists $t^{\star}>\ln\left(\sqrt{1+\frac{K}{\mu^{2}}}\right)$ such that (45) holds. ∎
@@ Appendix C Proof of the Main Results
We provide an upper bound for the moments of $(\widehat{Y}_{t}^{\text{EM}})_{t\in[0,T]}$ defined in (8) and an estimate for the one step error associated with $(\widehat{Y}_{t}^{\text{EM}})_{t\in[0,T]}$ which will be used in the proofs of Theorem 3.12 and Theorem 3.14 below.
Lemma C.1.
[7, Lemma 20] Let Assumptions 1 and 3.a hold. For any $p\in[2,4]$ and $t\in[0,T-\epsilon]$,
@@
$\begin{split}C_{\mathsf{EM},p}(t)&:=e^{t(3p-1-\frac{2}{p}+2^{2p-1}\mathsf{K}^{p}_{\text{Total}}(1+T^{\alpha p}))}\\ &\quad\times\left(\mathbb{E}\left[|\widehat{Y}_{0}^{\text{EM}}|^{p}\right]+2^{3p-2}\mathsf{K}^{p}_{\text{Total}}t(1+\mathbb{E}[|\hat{\theta}|^{p}])(1+T^{\alpha p})+\frac{2}{p}(pd+p(p-2))^{\frac{p}{2}}t\right).\end{split}$
Lemma C.2.
[7, Lemma 21] Let Assumptions 1 and 3.a hold. For any $p\in[2,4]$ and $t\in[0,T-\epsilon]$,
@@
and $C_{\mathsf{EM},p}$ is defined in Lemma C.1.
Proof of Theorem 3.12.
We derive the non-asymptotic estimate for $W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})$ using the splitting
$\begin{split}W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})&\leq W_{2}(\pi_{\mathsf{D}},\mathscr{L}(Y_{t_{K}}))+W_{2}(\mathscr{L}(Y_{t_{K}}),\mathscr{L}(\widetilde{Y}_{t_{K}}))\\ &\quad+W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K}}),\mathscr{L}(Y_{t_{K}}^{\text{aux}}))+W_{2}(\mathscr{L}(Y_{t_{K}}^{\text{aux}}),\mathscr{L}(Y_{K}^{\text{EM}})).\end{split}$ (49)
We provide upper bounds on the error made by the early stopping, i.e. $W_{2}(\pi_{\mathsf{D}},\mathscr{L}(Y_{t_{K}}))$, the error made by approximating the initial condition of the backward process $Y_{0}\sim\mathscr{L}(X_{T})$ with $\widetilde{Y}_{0}\sim\pi_{\infty}$, i.e. $W_{2}(\mathscr{L}(Y_{t_{K}}),\mathscr{L}(\widetilde{Y}_{t_{K}}))$, the error made by approximating the score function with $s$, i.e. $W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K}}),\mathscr{L}(Y_{t_{K}}^{\text{aux}}))$, and the discretisation error, i.e. $W_{2}(\mathscr{L}(Y_{t_{K}}^{\text{aux}}),\mathscr{L}(Y_{K}^{\text{EM}}))$, separately.
Upper bound on $W_{2}(\pi_{\mathsf{D}},\mathscr{L}(Y_{t_{K}})).$
The proof of this bound is the same as in the same bound provided in the [7, Proof of Theorem 10] using the representation of the OU process
[…]
$\begin{split}W_{2}(\pi_{\mathsf{D}},\mathscr{L}(Y_{t_{K}}))&\leq 2\sqrt{\epsilon}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d}),\end{split}$ (51)
where $t_{K}=T-\epsilon$.
Upper bound on $W_{2}(\mathscr{L}(Y_{t_{K}}),\mathscr{L}(\widetilde{Y}_{t_{K}})).$
@@
By integrating and taking on both sides in (52), we have
$\begin{split}\mathbb{E}\left[|Y_{t_{K}}-\widetilde{Y}_{t_{K}}|^{2}\right]&=\mathbb{E}\left[|Y_{0}-\widetilde{Y}_{0}|^{2}\right]+\int_{0}^{t_{K}}2\mathbb{E}\left[|Y_{t}-\widetilde{Y}_{t}|^{2}\right]\ \mathrm{d}t\\ &\quad+\int_{0}^{t_{K}}4\mathbb{E}\left[\langle Y_{t}-\widetilde{Y}_{t},\nabla\log p_{T-t}(Y_{t})-\nabla\log p_{T-t}(\widetilde{Y}_{t})\rangle\right]\ \text{d}t.\end{split}$ (53)
[…]
$\begin{split}&\mathbb{E}\left[|Y_{t_{K}}-\widetilde{Y}_{t_{K}}|^{2}\right]\\ &\leq\mathbb{E}\left[|Y_{0}-\widetilde{Y}_{0}|^{2}\right]+2\int_{0}^{t_{K}}\mathbb{E}\left[|Y_{t}-\widetilde{Y}_{t}|^{2}\right]\ \mathrm{d}t-4\int_{0}^{t_{K}}\beta^{\text{OS}}_{T-t}\mathbb{E}\left[|Y_{t}-\widetilde{Y}_{t}|^{2}\right]\text{d}t\\ &\leq\mathbb{E}[|Y_{0}-\widetilde{Y}_{0}|^{2}]e^{2[t_{K}-2\int_{0}^{t_{K}}\beta^{\text{OS}}_{T-t}\ \mathrm{d}t]}\\ &=\mathbb{E}[|m_{T}X_{0}+(\sigma_{T}-1)\widetilde{Y}_{0}|^{2}]e^{2[t_{K}-2\int_{0}^{t_{K}}\beta^{\text{OS}}_{T-t}\ \mathrm{d}t]}\\ &\leq 2\left(\mathbb{E}[|X_{0}|^{2}]+d\right)\ e^{2[t_{K}-2\int_{0}^{t_{K}}\beta^{\text{OS}}_{T-t}\ \mathrm{d}t]-2T}.\end{split}$ (54)
Using (54), Remark 3.9, and and $t_{K}=T-\epsilon$, we have
$\begin{split}W_{2}(\mathscr{L}(Y_{t_{K}}),\mathscr{L}(\widetilde{Y}_{t_{K}}))&\leq\sqrt{\mathbb{E}[|Y_{t_{K}}-\widetilde{Y}_{t_{K}}|^{2}]}\\ &\leq\sqrt{2}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d})e^{-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t-\epsilon}.\end{split}$ (55)
Upper bound on $W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K}}),\mathscr{L}(Y_{t_{K}}^{\text{aux}})).$
@@
By integrating and taking the expectation on both sides in (56), using Corollary 3.8, Young’s inequality with $\zeta\in(0,1)$ and Assumption 4, we have
$\begin{split}\mathbb{E}[|\widetilde{Y}_{T-\epsilon}-Y_{T-\epsilon}^{\text{aux}}|^{2}]&=2\int_{0}^{T-\epsilon}\mathbb{E}[|\widetilde{Y}_{s}-Y^{\text{aux}}_{s}|^{2}]\ \text{d}s\\ &\quad+4\int_{0}^{T-\epsilon}\mathbb{E}[\langle\widetilde{Y}_{s}-Y^{\text{aux}}_{s},\nabla\log p_{T-s}(\widetilde{Y}_{s})-\nabla\log p_{T-s}(Y_{s}^{\text{aux}})\rangle]\ \text{d}s\\ &\quad+4\int_{0}^{T-\epsilon}\mathbb{E}[\langle\widetilde{Y}_{s}-Y^{\text{aux}}_{s},\nabla\log p_{T-s}(Y_{s}^{\text{aux}})-s(T-s,\hat{\theta},Y_{s}^{\text{aux}})\rangle]\ \text{d}s\\ &\leq\int_{0}^{T-\epsilon}2(1+\zeta)\ \mathbb{E}[|\widetilde{Y}_{s}-Y^{\text{aux}}_{s}|^{2}]\ \text{d}s\\ &\quad-4\int_{0}^{t_{K}}\beta^{\text{OS}}_{T-s}\mathbb{E}\left[|\widetilde{Y}_{s}-Y^{\text{aux}}_{s}|^{2}\right]\text{d}t+2\zeta^{-1}\varepsilon_{\text{SN}}\\ &\leq 2e^{2(1+\zeta)(T-\epsilon)-4\int_{0}^{t_{K}}\beta^{\text{OS}}_{T-t}\ \mathrm{d}t}\ \zeta^{-1}\varepsilon_{\text{SN}}.\end{split}$ (57)
Using (57), Remark 3.9, and $t_{K}=T-\epsilon$, we have
$\begin{split}W_{2}(\mathscr{L}(\widetilde{Y}_{t_{K}}),\mathscr{L}(Y^{\text{aux}}_{t_{K}}))&\leq\sqrt{\mathbb{E}[|\widetilde{Y}_{t_{K}}-Y_{t_{K}}^{\text{aux}}|^{2}]}\\ &\leq\sqrt{2\zeta^{-1}}e^{(1+\zeta)(T-\epsilon)-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t}\sqrt{\varepsilon_{\text{SN}}}.\end{split}$ (58)
Upper bound on $W_{2}(\mathscr{L}(Y_{t_{K}}^{\text{aux}}),\mathscr{L}(\widehat{Y}_{t}^{\text{EM}})).$
@@
Integrating and taking the expectation on both sides in (59), using Young’s inequality for $\zeta\in(0,1)$, Cauchy Schwarz inequality, Assumption 3.a, Lemma C.2, and Remark 3.1, we have
$\begin{split}\mathbb{E}\left[|Y_{T-\epsilon}^{\text{aux}}-\widehat{Y}_{T-\epsilon}^{\text{EM}}|^{2}\right]&\leq(2+3\zeta)\int_{0}^{T-\epsilon}\mathbb{E}[|Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}}|^{2}]\ \text{d}t+\zeta^{-1}\int_{0}^{T-\epsilon}\mathbb{E}[|\widehat{Y}_{t}^{\text{EM}}-\widehat{Y}^{\text{EM}}_{\lfloor t/\gamma\rfloor\gamma}|^{2}]\ \text{d}t\\ &\quad+4\mathsf{K}_{3}(1+2T^{\alpha})\int_{0}^{T-\epsilon}\mathbb{E}[|Y_{t}^{\text{aux}}-\widehat{Y}_{t}^{\text{EM}}|^{2}]\mathrm{d}t\\ &\quad+2\zeta^{-1}\int_{0}^{T-\epsilon}\mathbb{E}[|s(T-t,\hat{\theta},\widehat{Y}^{\text{EM}}_{t})-s(T-\lfloor t/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor t/\gamma\rfloor\gamma}|^{2}]\mathrm{d}t\end{split}$

### v2

@@ Acknowledgments.
This work was supported by Innovate UK [grant number 10081810]. This work has received funding from the Ministry of Trade, Industry and Energy (MOTIE) and Korea Institute for Advancement of Technology (KIAT) through the International Cooperative R&D program (No.P0025828). This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program.
Appendix
@@ Appendix A Regularity of the Score Function
Proposition 23.
(Conforti et al., 2025, Proposition 3.1) Let $\pi_{\mathsf{D}}$ be absolutely continuous with respect to the Lebesgue measure, and denote its density by $p_{0}$. The map defined in 36 is positive and solution of the following Fokker–Planck equation on $(0,T]\times\mathbb{R}^{d}$:
@@ Appendix B Further Details on Assumption 2 and Weak Convexity of the Data Distribution
Proof of Proposition 13.
We begin by considering that $\pi_{\mathsf{D}}$ satisfies Assumption 2. Recall that $f_{L}$ is defined as in 13. Note that $r\mapsto r^{-1}f_{L}(r)$ is non-increasing on $(0,\infty)$ and $f_{L}^{{}^{\prime}}(0)=L>r^{-1}f_{L}(r)$ for $r\in(0,R]$. We look for $L>0$ satisfying
@@
so as $L>K+\mu$. Note that $\tanh(x)\leq x$ for all $x\geq 0$ . Therefore, if we choose $x=\frac{\sqrt{K+\mu}}{2}R$, then
@@
Using 39 and $\lim_{x\uparrow\infty}x\tanh(x)=\infty$, we deduce that there exists $x^{\star}>0$ such that
@@
with $x^{\star}>\frac{\sqrt{K+\mu}}{2}R$, since $x\mapsto x\tanh(x)$ is non-decreasing on $(0,\infty)$. By Assumption 2 and 37, we have
@@
where it is used that $r^{-1}f_{L}(r)>0$ for all $r>0$. This proves the first part of the statement in Proposition 13, i.e. the lower bound 16.
Conversely, assume that $U$ is weakly convex as in Definition 10 with lower bound 16 for some known constants $\mu$ and $L>0$. We look for $R$ such that
@@
where it is used that $r^{-1}f_{L}(r)$ is decreasing on $(0,\infty)$. Let $\widetilde{\mu}:=\mu-R^{-1}f_{L}(R)$, so 42 becomes $\kappa_{U}(r)\geq\widetilde{\mu}>0$, for all $r>R$. One notes that
@@
If $\mu>L$, 43 is satisfied for all $R>0$. If $\mu\leq L$, 43 holds for $R\geq R_{0}$, where $R_{0}$ is the unique solution to
@@
Since $\frac{\tanh(z)}{z}$ monotonically decreases from $1$ to 0 as $z$ increases, a unique $z_{0}>0$ solving 45 exists for $\mu<L$. Therefore, 42 is satisfied for $R\geq R_{0}=\frac{2z_{0}}{L^{1/2}}$. This proves that $U$ is $\widetilde{\mu}-$strongly convex at infinity, and therefore 17. Using the assumption that $U$ is weakly convex as in Definition 10, one obtains that
@@
We distinguish two cases for the lower bound in 46. If $\mu>L$, then $\kappa_{U}(r)\geq-K$ for $r\leq R$ for all $R>0$ and $K\geq 0$. If $\mu\leq L$, then, by setting $K=L-\mu$ in 45, we have $\kappa_{U}(r)\geq-K$ for $r\leq R$ for all $R>0$. This proves that $U$ is $K$-semiconvex, and therefore 18. This concludes the proof for the second part of the statement in Proposition 13.
Proof of Proposition 16.
@@
Note that 48 is satisfied for all $x>0$ when $K=0$. In addition, we have
@@
By 50, the function $g$ in 48 has a minimum at $\frac{K}{\mu^{2}}$ and
@@
for all $K,\mu>0$. By 49 and 50, there exists $x>\frac{K}{\mu^{2}}$ such that 48 is strictly positive. Therefore, there exists $t^{\star}>\ln\left(\sqrt{1+\frac{K}{\mu^{2}}}\right)$ such that 47 holds. ∎
@@ Appendix C Proof of the Main Results
In this section, we present the proofs of Theorem 19 and Theorem 21. We begin by recalling an upper bound on the moments of the process $(\widehat{Y}_{t}^{\text{EM}})_{t\in[0,T-\epsilon]}$ defined in 7, along with an estimate for its one-step discretization error. These results will be instrumental in the subsequent proofs.
Lemma 24.
(Bruno et al., 2025, Lemma 20) Let Assumptions 1 and 3.a hold, and suppose that $\mathbb{E}[|\hat{\theta}|^{p}]<\infty$ for any $p\in[2,4]$. Then, for any $t\in[0,T-\epsilon]$,
@@
$\begin{split}C_{\mathsf{EM},p}(t)&:=e^{t(3p-1-\frac{2}{p}+2^{2p-1}\mathsf{K}^{p}_{\text{Total}}(1+T^{\alpha p}))}\\ &\quad\times\left(\mathbb{E}\left[|\widehat{Y}_{0}^{\text{EM}}|^{p}\right]+2^{3p-2}\mathsf{K}^{p}_{\text{Total}}t(1+\mathbb{E}[|\hat{\theta}|^{p}])(1+T^{\alpha p})+\frac{2}{p}(pd+p(p-2))^{\frac{p}{2}}t\right),\end{split}$
and $\mathsf{K}_{\text{Total}}$ is defined in Remark 8.
Lemma 25.
(Bruno et al., 2025, Lemma 21) Let Assumptions 1 and 3.a hold, and suppose that $\mathbb{E}[|\hat{\theta}|^{p}]<\infty$ for any $p\in[2,4]$. Then, for any $t\in[0,T-\epsilon]$,
@@
$C_{\mathsf{EM},p}$ and $\mathsf{K}_{\text{Total}}$ are defined in Lemma 24 and in Remark 8, respectively.
Proof of Theorem 19.
We derive the non-asymptotic estimate for $W_{2}(\mathcal{L}(Y_{J}^{\text{EM}}),\pi_{\mathsf{D}})$ using the splitting
$\begin{split}W_{2}(\mathcal{L}(Y_{J}^{\text{EM}}),\pi_{\mathsf{D}})&\leq W_{2}(\pi_{\mathsf{D}},\mathcal{L}(Y_{t_{J}}))+W_{2}(\mathcal{L}(Y_{t_{J}}),\mathcal{L}(\widetilde{Y}_{t_{J}}))\\ &\quad+W_{2}(\mathcal{L}(\widetilde{Y}_{t_{J}}),\mathcal{L}(Y_{t_{J}}^{\text{aux}}))+W_{2}(\mathcal{L}(Y_{t_{J}}^{\text{aux}}),\mathcal{L}(Y_{J}^{\text{EM}})).\end{split}$ (51)
We provide upper bounds on the error made by the early stopping, i.e. $W_{2}(\pi_{\mathsf{D}},\mathcal{L}(Y_{t_{J}}))$, the error made by approximating the initial condition of the backward process $Y_{0}\sim\mathcal{L}(X_{T})$ with $\widetilde{Y}_{0}\sim\pi_{\infty}$, i.e. $W_{2}(\mathcal{L}(Y_{t_{J}}),\mathcal{L}(\widetilde{Y}_{t_{J}}))$, the error made by approximating the score function with $s$, i.e. $W_{2}(\mathcal{L}(\widetilde{Y}_{t_{J}}),\mathcal{L}(Y_{t_{J}}^{\text{aux}}))$, and the discretisation error, i.e. $W_{2}(\mathcal{L}(Y_{t_{J}}^{\text{aux}}),\mathcal{L}(Y_{J}^{\text{EM}}))$, separately.
Upper bound on $W_{2}(\pi_{\mathsf{D}},\mathcal{L}(Y_{t_{J}})).$
This bound can be established by following the same argument as in (Bruno et al., 2025, Proof of Theorem 10), which relies on the representation of the OU process
$X_{t}\overset{\text{a.s.}}{=}m_{t}X_{0}+\sigma_{t}Z_{t},\quad m_{t}=e^{-t},\quad\sigma_{t}^{2}=1-e^{-2t},\quad Z_{t}\sim\mathcal{N}(0,I_{d}),$ (52)
where $\overset{\text{a.s.}}{=}$ denotes almost sure equality. Therefore, we have
$\begin{split}W_{2}(\pi_{\mathsf{D}},\mathcal{L}(Y_{t_{J}}))&\leq 2\sqrt{\epsilon}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d}),\end{split}$ (53)
where $t_{J}=T-\epsilon$.
Upper bound on $W_{2}(\mathcal{L}(Y_{t_{J}}),\mathcal{L}(\widetilde{Y}_{t_{J}})).$
@@
By integrating and taking on both sides in 54, we have
$\begin{split}\mathbb{E}\left[|Y_{t_{J}}-\widetilde{Y}_{t_{J}}|^{2}\right]&=\mathbb{E}\left[|Y_{0}-\widetilde{Y}_{0}|^{2}\right]+\int_{0}^{t_{J}}2\mathbb{E}\left[|Y_{t}-\widetilde{Y}_{t}|^{2}\right]\ \mathrm{d}t\\ &\quad+\int_{0}^{t_{J}}4\mathbb{E}\left[\langle Y_{t}-\widetilde{Y}_{t},\nabla\log p_{T-t}(Y_{t})-\nabla\log p_{T-t}(\widetilde{Y}_{t})\rangle\right]\ \text{d}t.\end{split}$ (55)
By integrating, taking expectations on both sides in 55, using Corollary 14, the representation 52 with $Z_{T}\overset{\text{d}}{=}\widetilde{Y}_{0}$ (where $\overset{\text{d}}{=}$ denotes equality in distribution), the inequality $1-\sigma_{t}\leq m_{t}$ for any $t\in[0,T]$, we have
$\begin{split}&\mathbb{E}\left[|Y_{t_{J}}-\widetilde{Y}_{t_{J}}|^{2}\right]\\ &\leq\mathbb{E}\left[|Y_{0}-\widetilde{Y}_{0}|^{2}\right]+2\int_{0}^{t_{J}}\mathbb{E}\left[|Y_{t}-\widetilde{Y}_{t}|^{2}\right]\ \mathrm{d}t-4\int_{0}^{t_{J}}\beta^{\text{OS}}_{T-t}\mathbb{E}\left[|Y_{t}-\widetilde{Y}_{t}|^{2}\right]\text{d}t\\ &\leq\mathbb{E}[|Y_{0}-\widetilde{Y}_{0}|^{2}]e^{2[t_{J}-2\int_{0}^{t_{J}}\beta^{\text{OS}}_{T-t}\ \mathrm{d}t]}\\ &=\mathbb{E}[|m_{T}X_{0}+(\sigma_{T}-1)\widetilde{Y}_{0}|^{2}]e^{2[t_{J}-2\int_{0}^{t_{J}}\beta^{\text{OS}}_{T-t}\ \mathrm{d}t]}\\ &\leq 2\left(\mathbb{E}[|X_{0}|^{2}]+d\right)\ e^{2[t_{J}-2\int_{0}^{t_{J}}\beta^{\text{OS}}_{T-t}\ \mathrm{d}t]-2T}.\end{split}$ (56)
Using 56, Remark 15, and and $t_{J}=T-\epsilon$, we have
$\begin{split}W_{2}(\mathcal{L}(Y_{t_{J}}),\mathcal{L}(\widetilde{Y}_{t_{J}}))&\leq\sqrt{\mathbb{E}[|Y_{t_{J}}-\widetilde{Y}_{t_{J}}|^{2}]}\\ &\leq\sqrt{2}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{d})e^{-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t-\epsilon}.\end{split}$ (57)
Upper bound on $W_{2}(\mathcal{L}(\widetilde{Y}_{t_{J}}),\mathcal{L}(Y_{t_{J}}^{\text{aux}})).$
@@
By integrating and taking the expectation on both sides in 58, using Corollary 14, Young’s inequality with $\zeta\in(0,1)$ and Assumption 4, we have
$\begin{split}\mathbb{E}[|\widetilde{Y}_{T-\epsilon}-Y_{T-\epsilon}^{\text{aux}}|^{2}]&=2\int_{0}^{T-\epsilon}\mathbb{E}[|\widetilde{Y}_{s}-Y^{\text{aux}}_{s}|^{2}]\ \text{d}s\\ &\quad+4\int_{0}^{T-\epsilon}\mathbb{E}[\langle\widetilde{Y}_{s}-Y^{\text{aux}}_{s},\nabla\log p_{T-s}(\widetilde{Y}_{s})-\nabla\log p_{T-s}(Y_{s}^{\text{aux}})\rangle]\ \text{d}s\\ &\quad+4\int_{0}^{T-\epsilon}\mathbb{E}[\langle\widetilde{Y}_{s}-Y^{\text{aux}}_{s},\nabla\log p_{T-s}(Y_{s}^{\text{aux}})-s(T-s,\hat{\theta},Y_{s}^{\text{aux}})\rangle]\ \text{d}s\\ &\leq\int_{0}^{T-\epsilon}2(1+\zeta)\ \mathbb{E}[|\widetilde{Y}_{s}-Y^{\text{aux}}_{s}|^{2}]\ \text{d}s\\ &\quad-4\int_{0}^{t_{J}}\beta^{\text{OS}}_{T-s}\mathbb{E}\left[|\widetilde{Y}_{s}-Y^{\text{aux}}_{s}|^{2}\right]\text{d}t+2\zeta^{-1}\varepsilon_{\text{SN}}\\ &\leq 2e^{2(1+\zeta)(T-\epsilon)-4\int_{0}^{t_{J}}\beta^{\text{OS}}_{T-t}\ \mathrm{d}t}\ \zeta^{-1}\varepsilon_{\text{SN}}.\end{split}$ (59)
Using 59, Remark 15, and $t_{J}=T-\epsilon$, we have
$\begin{split}W_{2}(\mathcal{L}(\widetilde{Y}_{t_{J}}),\mathcal{L}(Y^{\text{aux}}_{t_{J}}))&\leq\sqrt{\mathbb{E}[|\widetilde{Y}_{t_{J}}-Y_{t_{J}}^{\text{aux}}|^{2}]}\\ &\leq\sqrt{2\zeta^{-1}}e^{(1+\zeta)(T-\epsilon)-2\int_{\epsilon}^{T}\beta^{\text{OS},K,\mu}_{t}\ \mathrm{d}t}\sqrt{\varepsilon_{\text{SN}}}.\end{split}$ (60)
Upper bound on $W_{2}(\mathcal{L}(Y_{t_{J}}^{\text{aux}}),\mathcal{L}(\widehat{Y}_{t}^{\text{EM}})).$
@@
Integrating and taking the expectation on both sides in 61, using Young’s inequality for $\zeta\in(0,1)$, Cauchy Schwarz inequality, Assumption 3.a, Lemma 25, and Remark 1, we have